Two surveyors with two-way radios leave the same point at 9:00 A.M., one walking due south at 4 mi/hr. How long can they communicate with one another if each radio has a maximum range of 2 miles?
Answer by stanbon(75887)

Two surveyors with two-way radios leave the same point at 9:00 A.M., one walking due south at 4 mi/hr. How long can they communicate with one another if each radio has a maximum range of 2 miles?
---------------------
If only one of them moves away from the point the following applies:
Walk Data:
distance = 2 miles ; rate = 4 mph; time = d/r = 2/4 = 1/2 hr
---------------
They can communicate for 1/2 hour.
Cheers,
Stan H.